FREQUENCY ANALYSIS OF THE 10 μm REGION OF THE ETHYLENE SPECTRUM USING THE D_{2h} TOP DATA SYSTEM.

M.-T. BOURGEOIS, M. ROTGER, Groupe de Spectrométrie Moléculaire et Atmosphérique, CNRS UMR 7331, Université de Reims Champagne-Ardenne, Moulin de la Housse, B.P. 1039, Cases 16-17, F-51687 Reims Cedex, France; M. TUDORIE, J. VANDER AUWERA, Service de Chimie Quantique et Photophysique, C.P. 160/09, Université Libre de Bruxelles, 50, avenue F. D. Roosevelt, B-1050 Brussels, Belgium; \underline{V. BOUDON}, Laboratoire Interdisciplinaire Carnot de Bourgogne, CNRS UMR 6303, Université de Bourgogne, 9, Avenue Alain Savary, BP 47870, F-21078 Dijon Cedex, France.

New high resolution IR spectra of the ν_{10}/ν_{7}/ν_{4} region, lying from 600 cm^{−1} to 1200 cm^{−1} have been recorded in Brussels. They are used simultaneously with a previous spectrum of the ν_{12} band, also recorded in Brussels, to perform the analysis of the ν_{10}/ν_{7}/ν_{4}/ν_{12} region, lying from 600 cm^{−1} to 1500 cm^{−1}. Following our work devoted to the ν_{12} band considered as isolated, we performed a frequency re-analysis of the ν_{10}/ν_{7}/ν_{4}/ν_{12} infrared tetrad in the 600 − 1500 cm^{−1} region, thanks to the tensorial formalism developed in Dijon for X_{2}Y_{4} asymmetric-top molecules^{a} and the XTDS software package^{b}. Including the strong Coriolis interaction affecting the upper vibrational levels 10^{1}, 7^{1}, 4^{1} and 12^{1}, a total of 11727 lines have been assigned and fitted as a tetrad with a global root mean square deviation of 0.338 × 10^{−3} cm^{−1}.

aW. Raballand, M. Rotger, V. Boudon and M. Loëte, J. Mol. Spectrosc., 217 239–248 (2003).

bCh. Wenger, V. Boudon, M. Rotger, M. Sanzharov and J.-P. Champion, J. Mol. Spectrosc., 251 102–113 (2008).